If the difference of the roots of 2xx-6x+c=0 is 5, find the roots and the value of c.
Answer by TimothyLamb(4379)

2xx - 6x + c = 0
---
complete the square:
---
(2xx - 6x + c = 0)/2
xx - 3x + c/2 = 0
xx - 3x + (-3/2)^2 - (-3/2)^2 = -c/2
xx - 3x + (-3/2)^2 = -c/2 + (-3/2)^2
(x - 3/2)^2 = -c/2 + 9/4
---
solve for x:
---
sqrt( (x - 3/2)^2 ) = sqrt( -c/2 + 9/4 )
x - 3/2 = sqrt( -c/2 + 9/4 )
x = sqrt( -c/2 + 9/4 ) + 3/2
x = sqrt( -2c/4 + 9/4 ) + 3/2
x = sqrt( (9 - 2c)/4 ) + 3/2
x = sqrt( 9 - 2c )/2 + 3/2
---
roots are:
x = 3/2 + sqrt( 9 - 2c )/2
x = 3/2 - sqrt( 9 - 2c )/2
---
difference of roots is 5:
3/2 + sqrt( 9 - 2c )/2 - ( 3/2 - sqrt( 9 - 2c )/2 ) = 5
3/2 + sqrt( 9 - 2c )/2 - 3/2 + sqrt( 9 - 2c )/2 = 5
sqrt( 9 - 2c )/2 + sqrt( 9 - 2c )/2 = 5
sqrt( 9 - 2c ) = 5
( sqrt( 9 - 2c ) )^2 = 5*5
9 - 2c = 25
2c = 9 - 25
c = -16/2
c = -8
---
roots:
x = 3/2 + sqrt( 9 - 2*(-8) )/2
x = 3/2 - sqrt( 9 - 2*(-8) )/2
---
x = 4
x = -1
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---